if you transfer 0.1 ml of a sample into a 99.9ml saline blank, what is the dilution?
Answer #1

For finding dilution ratio, simple equation can be given as

=ml of sample/(ml of sample+ml of diluent)

=(0.1/(0.1+99.9))

=0.1/100

=1/1000

So, this result gives us dilution, which is 1/1000